You deposit $3000 into a bank account that compounds interest quarterly. If the APR is 2.5%, how much will you have in 5 years? Round to the nearest dollar, no commas. *


1
Expert's answer
2020-08-12T18:26:29-0400

A=P(1 + r/n)nt

Where;

A=the new balance

P= principle amount ($3000).

r= interest rate in decimal form (0.025).

n= the number of times it is compounded (4).

t= the time 5 years.

A=P(1 + r/n)nt

A=3000(1+ 0.025/4)(5)(4)

=3000(1+ 0.00625)20

=3000(1.00625)20

=3398.1232

=$ 3398